OpenStudy (x3_drummerchick):

Verify the trig function: can someone explain to me how my professor got from step 1 to 2? will give medals!

OpenStudy (x3_drummerchick):

i attached a photo of the problem

OpenStudy (x3_drummerchick):

@freckles @Nnesha

Nnesha (nnesha):

so he is working on the right side

OpenStudy (x3_drummerchick):

yes

Nnesha (nnesha):

i'm sure you're familiar with the special identity \[\huge\rm \sin^2x+\cos^2x=1\] solve this identity for sin^2x

OpenStudy (x3_drummerchick):

1-cos^2x=sin^2x

Nnesha (nnesha):

right so now we can replace sin^2x with 1-cos^2 \[\huge\rm \tan^2x \color{reD}{sin^2x} \] \[\huge\rm \tan^2x \color{reD}{(1-cos^2x)} \]

OpenStudy (x3_drummerchick):

right

OpenStudy (x3_drummerchick):

but how did they go from that ^ to tan^2-(sin^2/cos^2)?

Nnesha (nnesha):

that's what i'm trying to understand .... ,-,

Nnesha (nnesha):

can you take screenshot of next step

OpenStudy (x3_drummerchick):

he said something about rewriting it as tan, but it does make sense

OpenStudy (x3_drummerchick):

doesnt *

Nnesha (nnesha):

it doesn't make sense i guess he is trying to rewrite tan^2 in terms of sin and cos

Nnesha (nnesha):

\[\huge\rm \tan^2x \color{reD}{(1-cos^2x)} \]\[\large\rm \frac{ \sin^2x }{ \cos^2x }(1-\cos^2x)\]

OpenStudy (x3_drummerchick):

im still confused :(

Nnesha (nnesha):

hm let me think

Nnesha (nnesha):

ahh i see so he distributed by sin^2/cos^2 \[\rm \frac{ \sin^2 x}{ \cos^2x }-\frac{\sin^2}{\cos^2x}*\cos^2x\] and then he changed sin^2/cos^2 to tan^2

OpenStudy (x3_drummerchick):

im sorry, could you rewrite it out from the start, im still confused

Nnesha (nnesha):

\[\huge\rm \tan^2x \color{reD}{(1-cos^2x)} \]\[\large\rm \color{blue}{ \frac{ \sin^2x }{ \cos^2x} }(1-\cos^2x)\] rewrite tan as sin/cos \[\rm \color{blue}{ \frac{ \sin^2 x}{ \cos^2x }}-\frac{\sin^2}{\cos^2x}*\cos^2x\] distribute parentheses by sin^2/cos^2 \[\rm \color{blue}{\tan^2x} -\frac{sin^2x}{cos^2x}*cos^2x\] rewrite sin^2/cos^2 as tan^2

OpenStudy (x3_drummerchick):

im still lost on this, im sorry

OpenStudy (x3_drummerchick):

ohhh wait i think i get ittt

OpenStudy (x3_drummerchick):

he distributed tan from the get go?

Nnesha (nnesha):

okay so i applied distributive property \[\large\rm \color{ReD}{a}(b+c)=\color{ReD}{a}*b+\color{ReD}{a}*c=\color{red}{a}b+\color{ReD}{a}c\] multiply both terms in the parentheses by outside term that's how  \[\rm \color{blue}{ \frac{ \sin^2 x}{ \cos^2x }}-\frac{\sin^2}{\cos^2x}*\cos^2x\]